If someone would please explain this question to me. I use this symbol ( / )for the absolute value.
/3-4/-/5-7/+/-2+6/=

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

/3-4/-/5-7/+/-2+6/=
--------
= |-1| - |-2| + |4|
------
= 1 - 2 + 4
---
= 3
==============
cheers,
Stan H.
================

Answer by edjones(8007)

Use the "|" sign which comes on when the Shift key is depressed and the backward slash "\" key is pressed at the same time.
|3-4|-|5-7|+|-2+6|
=|-1|-|-2|+|4|
=1-2+4
=3
.
Ed